Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 432     Entry time: 02 Nov 2023, 23:25
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 19 keV, ~0.9 mJ/pulse, stable delivery

Optical laser delivery

  •  no

Achievements/Observations

  • Tried many possibilities to fix the jitter issue. We covered everything we could think of, but in the end it was not possible as we could not avoid restricting ourselves from the 0.4s delay required for the ultrasound, making us then suffer from the inter-train jitter.
  • Returned to operational experiment setup, eventhough suffering from the jitter issue.

Issues

  • High jitter from output of transducer (Ch 2in oscilloscope) could not be solved.
  • Turbopumps went off, then we realized multiple power sockets were powerless (F7,8,9,10,11). We connected the turbos to another socket to resume operation, after which the previous sockets were also fixed. We must reconnect the turbos to the typical sockets sometime soon, best expectation to be comming Monday.
ELOG V3.1.4-7c3fd00